Find the nature and critical point of the system
dx/dt = (x)^2 +3xy+2(y)^2 and
dy/dt= 4-(x)^2
1
Expert's answer
2021-07-27T09:08:49-0400

The only critical point is the solution of the simultaneous equations


"(x)^2 +3xy+2(y)^2=0"

"4-(x)^2=0"

"x=-2: 4-6y+2y^2=0"

"y^2-3y+2=0"

"(y-1)(y-2)=0"

"Point(-2, 1), Point (-2, 2)"


"x=2: 4+6y+2y^2=0"

"y^2+3y+2=0"

"(y+1)(y+2)=0"

"Point(2, -2), Point (2, -1)"



Critical points of the system


"Point(-2, 1), Point (-2, 2),"

"Point(2, -2), Point (2, -1)"


The Jacobian matrix is


"J=\begin{pmatrix}
 \dfrac{\partial f}{\partial x} & \dfrac{\partial f}{\partial y} \\
\\
 \dfrac{\partial g}{\partial x} & \dfrac{\partial g}{\partial y}
\end{pmatrix}=\begin{pmatrix}
 2x+3y & 3x+4y \\
 -2x & 0
\end{pmatrix}"


At the "Point(-2, 1)"  we get the matrix "\begin{pmatrix}
 -1 & -2 \\
 4& 0
\end{pmatrix}" and so the two eigenvalues are 

"\dfrac{-1-i\sqrt{31}}{2}" and "\dfrac{-1+i\sqrt{31}}{2}."

As the eigenvalues are complex with negative real part, we get a spiral sink, which is an asymptotically stable equilibrium point.


At the "Point(-2, 2)"  we get the matrix "\begin{pmatrix}
 2 & 2 \\
 4 & 0
\end{pmatrix}" and so the two eigenvalues are "-2" and "4."

As the eigenvalues are real and opposite signs, we get a saddle, which is an unstable equilibrium point.


At the "Point(2, -2)"  we get the matrix "\begin{pmatrix}
 -2 & -2 \\
 -4 & 0
\end{pmatrix}" and so the two eigenvalues are "-4" and "2."

As the eigenvalues are real and opposite signs, we get a saddle, which is an unstable equilibrium point.


At the "Point(2, -1)"  we get the matrix "\begin{pmatrix}
 1 & 2 \\
 -4 & 0
\end{pmatrix}" and so the two eigenvalues are 

"\dfrac{1-i\sqrt{31}}{2}" and "\dfrac{1+i\sqrt{31}}{2}."

As the eigenvalues are complex with positive real part, we get a spiral source, which is an unstable equilibrium point.